# Duality and Axion Wormholes

### v1

@@ 1 Introduction
(The precise factor ${\mathrm{i}}f/3!h$ ensures that the stress tensor transforms properly, leading to an equivalence of the Einstein field equations coupled to $\phi$ or $B$. The overall sign is an arbitrary convention.) In Lorentz signature, a similar relation without the factor of $\mathrm{i}$} does largely, but not fully, account for the duality, as explained in section 2. In Euclidean signature, the formula requires an imaginary factor, and it does not make sense as a relationship between classical fields, since it is not consistent with Dirac quantization. Dirac quantization requires ${\mathrm{d}}\phi$ and $H$ to have real periods (indeed, periods that are integer multiples of $2\pi$), but the factor of $\mathrm{i}$} in (4) makes this impossible if (4) is regarded as a relation between classical fields. Actually, though it does not make sense in Euclidean signature as a transformation between classical fields, eqn. (4) is valid in Euclidean signature as a statement that holds inside quantum correlation functions. This will be explained in section 2.
@@
More specifically, in the case of the wormhole, since the three-form $H$ has a non-zero (and real!) flux when integrated on a cross-section of the wormhole throat, the relation (4) would require $\phi$ to change by an imaginary constant in going “through the wormhole” from one asymptotically flat region to the other. In general, in the gravitational path integral, one considers classical fields that are not real, but one does want to impose suitable reality conditions in asymptotic regions of spacetime. In the case of the $\phi$ field, one wants $\phi$ to equal a real constant at infinity; the value of this constant represents a choice of vacuum. If $\phi$ changes by an imaginary amount in going through the wormhole, it cannot be real at both ends. It would be highly problematical to give $\phi$ an imaginary value at infinity. To see this, consider elaborating the theory slightly by adding gauge fields with field strength $F$ and an axion coupling that in Euclidean signature is ${\mathrm{i}}\int\phi\,{\mathrm{Tr}}\,F\wedge F$ (in Lorentz signature there is no factor of $\mathrm{i}$} in this coupling). If $\phi$ has an imaginary expectation value, the gauge fields have an imaginary theta-angle, violating unitarity.
@@
Despite all this, in more complicated theories with one or more two-forms coupled to scalars, the relationship (4) and its generalizations is very useful; it has been used to efficiently analyze wormhole solutions of low energy effective actions derived from string theory AOP . In general, as shown in that paper, a very efficient way to find wormhole solutions of a theory with multiple scalars and two-forms is to convert all of the two-forms to imaginary scalars by relations such as (4), solve the equations of motion in the scalar language, and then convert the imaginary scalars back to two-forms via (4).
@@ 2.1 Dirac Quantization and Relation Between Couplings
The relation
@@
almost defines an on-shell transformation between a $B$-field with field strength $H={\mathrm{d}}B$ and a massless scalar $\phi$. This is most obvious if the transformation is written in terms of differential forms. In that language, eqn. (5) reads ${\mathrm{d}}\phi=\frac{f}{h}\star H$, where $\star$ is the Hodge star operator. It exchanges two equations satisfied by $\phi$, namely the classical equation of motion ${\mathrm{d}}\star{\mathrm{d}}\phi=0$ and the identity ${\mathrm{d}}^{2}\phi=0$, with two equations satisfied by $H$, namely the Bianchi identity ${\mathrm{d}}H=0$ and the classical equation of motion ${\mathrm{d}}\star H=0$. Indeed, if ${\mathrm{d}}\phi=\frac{f}{h}\star H$, then the identity ${\mathrm{d}}^{2}\phi=0$ is equivalent to the classical equation ${\mathrm{d}}\star H=0$, and the classical equation ${\mathrm{d}}\star{\mathrm{d}}\phi=0$ is equivalent to the Bianchi identity ${\mathrm{d}}H=0$.
The constant $f/3!h$ in eqn. (5) is chosen so that if $\phi$ and $H$ are related in this way, then the stress tensor of $\phi$ equals the stress tensor of $H$. (In Euclidean signature, this condition requires an extra factor of $\pm{\mathrm{i}}$, as in eqn. (4).) It is also true that the canonical commutation relations of $\phi$ are mapped to those of $B$ by the relation (5). Thus, (5) almost suffices to characterize a quantum duality between the scalar and the two-form. But eqn. (5) is not a complete relationship between $\phi$ and $B$, because it omits important information about zero-modes on the two sides. If $H$ is given, then $\phi$ is only determined up to an additive constant. And if $\phi$ is given, then eqn. (5) determines $H$, but this does not quite suffice to determine $B$ up to a gauge transformation, because there can be $B$-fields that are not pure gauge, but have $H=0$ (for example, a $B$-field on a two-manifold $\Sigma_{2}$ with $\int_{\Sigma_{2}}B\not=0$). To extend the duality to encompass the zero-modes, one has to impose Dirac quantization on both sides of the duality. Moreover, for the zero-modes, the duality cannot be expressed by a classical mapping like (5) between scalars and two-forms, but involves quantization.
@@ Appendix B Some More Details About the Path Integral
The ratio of the $B$-field path integral to the scalar path integral is then $f^{b_{2}-2b_{1}+2b_{0}}=f^{\chi}$, where $\chi=b_{0}-b_{1}+b_{2}-b_{3}+b_{4}=2b_{0}-2b_{1}+b_{2}$ is the Euler characteristic of $M$, or an $L^{2}$ version of the Euler characteristic if $M$ is not compact. This result was first obtained in [^11] That paper also contains a much more precise computation than we have explained here. Our discussion has not been precise enough to determine whether the anomalous factor is $f^{\chi}$ or $(cf)^{\chi}$ for some constant $c$. A similar anomalous factor in electric-magnetic duality in four dimensions was found in Witten . DMW . If $M$ is compact, the Euler characteristic of $M$ can be written as a curvature integral:
@@
that depends only on the background metric, we can restore the equivalence between the scalar and $B$-field path integrals. We should acknowledge, though, that from the point of view of the derivation in section 2.2, it is not very clear why this counterterm is needed. If $M$ is not compact, then the $L^{2}$ Euler characteristic of $M$ is given by the formula (102) plus a local boundary correction that is somewhat analogous to the Gibbons-Hawking-York boundary correction to the Einstein-Hilbert action. That boundary term must be included in the action.
@@
Acknowledgements I thank A. Heiderschee, V. Ivo, J. Maldacena, and especially G. Moore for comments and discussions. Research supported in part by NSF Grant PHY-2514611.

### v2

@@ 1 Introduction
(The precise factor ${\mathrm{i}}f/3!h$ ensures that the stress tensor transforms properly, leading to an equivalence of the Einstein field equations coupled to $\phi$ or $B$. The overall sign is an arbitrary convention.) In Lorentz signature, a similar relation without the factor of $\mathrm{i}$} does largely, but not fully, account for the duality, from a Hamiltonian point of view, as explained in section 2. This is not the case for path integrals. The relation (4) implies the classical equations of motion for both $\phi$ and $H$, and hence cannot be defined off-shell. So it does not make sense as a transformation of the fields in a path integral, which are generically off-shell. But eqn. (4) can be deduced from path integrals as a statement that holds inside quantum correlation functions. This will be explained in section 2.
Even on-shell, there is a problem with interpreting eqn. (4) as a relation between classical fields. Dirac quantization requires ${\mathrm{d}}\phi$ and $H$ to have real periods (indeed, periods that are integer multiples of $2\pi$), but the factor of $\mathrm{i}$} in (4) makes this impossible if (4) is regarded as a relation between classical fields. Even in Lorentz signature, where there is no factor off $\mathrm{i}$} in the formula, the factor of $f/h$ means that as a relation between classical fields, eqn. (4) is generically incompatible with Dirac quantization.
@@
In the particular case of the wormhole, since the three-form $H$ has a non-zero (and real!) flux when integrated on a cross-section of the wormhole throat, the relation (4) would require $\phi$ to change by an imaginary constant in going “through the wormhole” from one asymptotically flat region to the other. In general, in the gravitational path integral, one considers classical fields that are not real, but one does want to impose suitable reality conditions in asymptotic regions of spacetime. In the case of the $\phi$ field, one wants $\phi$ to equal a real constant at infinity; the value of this constant represents a choice of vacuum. If $\phi$ changes by an imaginary amount in going through the wormhole, it cannot be real at both ends, and hence the solution obtained by applying the transformation (4) to the wormhole solution of GS does not satisfy the boundary conditions that one would want in wormhole physics.
Indeed, it would be highly problematical to give $\phi$ an imaginary value at infinity. To see this, consider elaborating the theory slightly by adding gauge fields with field strength $F$ and an axion coupling that in Euclidean signature is ${\mathrm{i}}\int\phi\,{\mathrm{Tr}}\,F\wedge F$ (in Lorentz signature there is no factor of $\mathrm{i}$} in this coupling). If $\phi$ has an imaginary expectation value, the gauge fields have an imaginary theta-angle, violating unitarity.
@@
Despite all this, in more complicated theories with one or more two-forms coupled to scalars, the relationship (4) and its generalizations are very useful; they have been used to efficiently analyze wormhole solutions of low energy effective actions derived from string theory AOP . In general, as shown in that paper, a very efficient way to find wormhole solutions of a theory with multiple scalars and two-forms is to convert all of the two-forms to imaginary scalars by relations such as (4), solve the equations of motion in the scalar language, and then convert the imaginary scalars back to two-forms via (4).
@@ 2.1 Dirac Quantization and Relation Between Couplings
The relation [^2] In Euclidean signature, the following relation requires an extra factor of $\pm{\mathrm{i}}$; see eqn. (4).
@@
almost defines an on-shell transformation between a $B$-field with field strength $H={\mathrm{d}}B$ and a massless scalar $\phi$. This is most obvious if the transformation is written in terms of differential forms. In that language, eqn. (5) reads ${\mathrm{d}}\phi=\frac{f}{h}\star H$, where $\star$ is the Hodge star operator. It exchanges two equations satisfied by $\phi$, namely the classical equation of motion ${\mathrm{d}}\star{\mathrm{d}}\phi=0$ and the identity ${\mathrm{d}}^{2}\phi=0$, with two equations satisfied by $H$, namely the Bianchi identity ${\mathrm{d}}H=0$ and the classical equation of motion ${\mathrm{d}}\star H=0$. Indeed, if ${\mathrm{d}}\phi=\frac{f}{h}\star H$, then the identity ${\mathrm{d}}^{2}\phi=0$ is equivalent to the classical equation ${\mathrm{d}}\star H=0$, and the classical equation ${\mathrm{d}}\star{\mathrm{d}}\phi=0$ is equivalent to the Bianchi identity ${\mathrm{d}}H=0$. The constant $f/3!h$ in eqn. (5) is chosen so that if $\phi$ and $H$ are related as in eqn. (5), then the stress tensor of $\phi$ equals the stress tensor of $H$.
Because it exchanges the equations of motion with the Bianchi identities ${\mathrm{d}}({\mathrm{d}}\phi)={\mathrm{d}}H=0$, the transformation (5) implies the equations of motion and cannot be defined off-shell. So it does not make sense as a transformation of the classical fields in a path integral, which are generically off-shell. (As we will see, it can be deduced from path integrals as a statement about quantum correlation functions; roughly, from the point of view of path integrals, this relation holds on the average.) In a Hamiltonian approach, one constructs a Hilbert space by quantizing the classical phase space of a theory. The classical phase space can be viewed as the space of classical solutions of a theory, so in a Hamiltonian approach, one never has to consider off-shell fields; one only has to consider on-shell fields and their canonical commutation relations. The relation (5) does map the canonical commutation relations of $\phi$ to those of $B$. So, in a Hamiltonian approach, (5) almost suffices to characterize a quantum duality between the scalar and the two-form. But eqn. (5) does not define a complete quantum duality in a Hamiltonian approach, because it omits important information about zero-modes on the two sides. If $H$ is given, then $\phi$ is only determined up to an additive constant. And if $\phi$ is given, then eqn. (5) determines $H$, but this does not quite suffice to determine $B$ up to a gauge transformation, because there can be $B$-fields that are not pure gauge, but have $H=0$ (for example, a $B$-field on a two-manifold $\Sigma_{2}$ with $\int_{\Sigma_{2}}B\not=0$). To extend the duality to encompass the zero-modes, one has to impose Dirac quantization on both sides of the duality, and impose a certain relation between $f$ and $h$. Moreover, for the zero-modes, the duality cannot be expressed by a classical mapping like (5) between scalars and two-forms, but involves quantization.
@@ Appendix B Some More Details About the Path Integral
The ratio of the $B$-field path integral to the scalar path integral is then $f^{b_{2}-2b_{1}+2b_{0}}=f^{\chi}$, where $\chi=b_{0}-b_{1}+b_{2}-b_{3}+b_{4}=2b_{0}-2b_{1}+b_{2}$ is the Euler characteristic of $M$, or an $L^{2}$ version of the Euler characteristic if $M$ is not compact. This result was first obtained in [^12] That paper also contains a much more precise computation than we have explained here, fixing an overall constant that we have not determined (is the anomaly precisely $f^{\chi}$ or should $f$ be multiplied by a constant?). A similar anomalous factor in electric-magnetic duality in four dimensions was found in Witten . In that reference, electric-magnetic duality was derived in a similar fashion to what was presented here in section 2.2, and the origin of the Euler characteristic anomaly was explained in that framework. DMW . If $M$ is compact, the Euler characteristic of $M$ can be written as a curvature integral:
@@
that depends only on the background metric, we can restore the equivalence between the scalar and $B$-field path integrals. If $M$ is not compact, then the $L^{2}$ Euler characteristic of $M$ is given by the formula (102) plus a local boundary correction that is somewhat analogous to the Gibbons-Hawking-York boundary correction to the Einstein-Hilbert action. That boundary term must be included in the action.
@@
Acknowledgements I thank N. Arkani-Hamed, A. Heiderschee, V. Ivo, J. Maldacena, and especially G. Moore for comments and discussions. Research supported in part by NSF Grant PHY-2514611.